OpenStudy (anonymous):

PLEASE HELP!!!!!! ]Choose the equation of the line passing through the point (3, -2) and parallel to y = 3x - 6.

OpenStudy (anonymous):

u get -1/3 because the product of slopes of perpendicular lines is -1 so x*3=-1 so x=-1/3

OpenStudy (anonymous):

so would the answer be y = 3x - 11

OpenStudy (amistre64):

all lines that are parallel to that one are related to: y = 3x we need to adjust the given point (3,-2) so that it is the origin (0,0) (3-3,-2+2) but we need to do this to not only this one point, but ALL the points (x,y) on the line (x-3,y+2) now we can use this in our related equation: (y+2) = 3(x-3) and simplify to your hearts content

OpenStudy (amistre64):

so yes; y = 3x - 11 is good